OpenStudy (anonymous):

It takes 158 s for an unknown gas to effuse through a porous wall and 84.0 s for the same volume of N2 gas to effuse at the same temperature and pressure. What is the molar mass of the unknown gas? No answers, I want a solution please.

OpenStudy (anonymous):

I don't know about this kind of problem. Can you enlighten me? Are we using the formula \[v _{1}^{2}\times M _{1}=v _{2}^{2}\times M _{2}\]

OpenStudy (anonymous):

I found the solution already. It's basically using Grahm's Law of Diffusion, which is not the same as effusion but to solve the problem, the same formula is used. As you can see, rate and the square root of the molar mass are inversely proportional. \[\frac{ r _{1} }{ r _{2}} = \sqrt{\frac{ M _{2} }{ M _{1} }}\]Time is also inversely proportional to the rate so the formula now becomes: \[\frac{ t _{1} }{ t _{2} } = \sqrt{\frac{ MM _{1} }{ MM _{2} }}\] Now plug the numbers in, and the problem is solved.

OpenStudy (anonymous):

Awesome.